OpenStudy (anonymous):

a jeweler has five rings, each weighing 18, made of an alloy of 10% silver and 90% gold. She decides to malt down the rings and add enough silver to reduce the gold content for 75%. How much silver should she add?

OpenStudy (anonymous):

@Isaiah.Feynman can you please help my friend?? thanks

OpenStudy (anonymous):

total weight is 5(18) = 90. And 10% are silver. Let x = amount of silver added, then 0.10(90) + x = 0.25(90+x)

OpenStudy (anonymous):

can you explain this a lil more for me @Isaiah.Feynman

OpenStudy (anonymous):

okay but the answer is 18 but im getting 22.5

OpenStudy (mathstudent55):

You don't specify the units, but I'll use grams. The total weight of all 5 rings is 5 * 18 g = 90 g. In 90 g of alloy, 90% is gold, that means 90% * 90 g = 81 g of gold. Since you want the alloy now to be 75% gold, solve the percentage problem below: 81 g is 75% of what? 81 g = 75% * x 81 g = 0.75x x = 108 g That means the total weight of the rings must now become 108 g. Since they already weigh 90 grams, 108 g - 90 g = 18 g He needs to add 18 grams of silver.

OpenStudy (anonymous):

gracias!

OpenStudy (mathstudent55):

de nada